# Precomputed Dominant Resource Fairness

### v1

@@ 5 Performance of PDRF
In the original DRF article Ghodsi et al. (2011) point out that each allocation operation takes $\mathcal{O}(log(n))$ time, since the selection of lowest allocated dominant share may be done with a binary heap. However, the number of allocation operations is determined by the ratio of the available resource reserves to the demand averages. More rigorously, the expected execution time of the algorithm is given by:
@@
where, $D$ and $R$ denote the set of demands and resources, respectively, and $\mu_{dr}$ is the demand average for resource $r$. This can lead to performance problems in cases where this ratio is relatively high.
The main contribution of PDRF is that in such cases it can reduce the time of execution dramatically, since it operates in linear time to the number of users and resources, i.e. $\mathcal{O}(m,n)$, and is not effected from this ratio. It takes $m\cdot n$ multiplications to find the fractional demands, $(m-1)\cdot n$ comparisons to find the $ds$ of each demand, $n-1$ comparisons to find $ds^{*}$, $n$ divisions for finding the ratio of $ds^{*}$ to each $ds$, $m\cdot n$ scaling of demands by the obtained ratios, $m\cdot(n-1)$ additions for summing up the scaled demands, $m$ divisions of resources by the demand sums, $m-1$ comparisons to find the minimum of the results of the divisions to decide the number of of cycle iterations, and $2\cdot m\cdot n$ operations for the final assignment of resources, all of which are linear time operations.
Although a final allocation round employing DRF may be needed to satisfy underallocated demands and reach Pareto Efficiency, it iterates over $47\%$ of the number of users, on average, and each of them once, since as shown in Section 4, that is the expected number of users that are short by $1$ task to the DRF allocation. Yet better, since all underallocations are by $1$ task, simply allocating $1$ task, starting from the least allocated user and proceeding in the ascending order until demands are exhausted may suffice, rather than a DRF iteration, which takes $2$ heap operations for user (one taking out and one putting back), and an assignment operation for updating the allocated dominant share value.
@@ 6.1 Higher Order Cycles
Without loss of generality, we may consider the following single resource scenario where the demands are $2$, $4$, and $10$, for users A, B and C, respectively, and the reserve is $35$ units. User C has the $ds^{*}$, so one cycle iteration consists of $1$ iteration for C, $2$ for B, $4$ for A, and it consumes $2\cdot 4+4\cdot 2+10\cdot 1=26$ units of the reserve. After one iteration of the cycle, $9$ units of resource is left and the cycle cannot take another turn. When we exclude User C, who has the $ds^{*}$, what we are left with is a secondary cycle, with now B having the new $ds^{*}$ and taking $1$ unit, and A taking $2$. The cycle takes one turn consuming 8 resources, and terminates.
@@ 6.2 Pareto Efficiency
The proof is based on the claim that "progressive filling for DRF is equivalent to the algorithm presented in Figure $1$" [^1] Since Figure $1$ is a scatter plot, we assume what is meant is Algorithm $1$. (Ghodsi et al., 2011, p. 13). Although it is indicated in the preceding paragraph that the progressive filling algorithm eliminates the saturated demands, no removal condition is given in Algorithm $1$, and neither it is mentioned in the sections and examples describing the algorithm.This might be a minor point eluding the authors’ attention or it might have been assumed that the selection process automatically excludes those demands, but the resulting algorithm as such, cannot achieve Pareto Efficiency in, though rare might they be, certain demand distributions and resource reserves.
At the moment we cannot provide an exact mathematical model for the construction of such configurations, but when the $ds^{*}$ is saturated, and there are more resources available for any other lower demand than its number of occurrences in the cycle, meaning though saturated, $ds^{*}$ will come to be the least allocated with the beginning of the next cycle, the algorithm terminates without allocating available resources to the other users.
An example we rather stumbled upon, which led to this analysis, consists of the same demand vectors in the original article, user A with the demand vector $\left\langle 1,4\right\rangle$ and user B with $\left\langle 3,1\right\rangle$, and the available capacities $\left\langle 59,19\right\rangle$. In such a setting, after 10 iterations, the algorithm halts allocating $\left\langle 2,8\right\rangle$ and $\left\langle 24,8\right\rangle$, respectively, and $\left\langle 33,3\right\rangle$ resources still available, which is sufficient for $3$ more allocations for user B, but since the least allocated dominant share belongs to A, and his demand is saturated, the algorithm cannot proceed. A short simulator script with this example can be found at the same github repository.
As in EDRF, this problem may easily be overcome by excluding the saturated demand from the demand set.
@@ 6.3 Weighting Policies
A weighting policy, assigns weights to users, and allocates the resources in proportion to these weights. Fairness, in turn, is defined over the allocations scaled by these weights. In Ghodsi et al. (2011), the question of weighting is addressed briefly, pointing to the fact that it suffices for such a generalisation to assign a weight vector to each user, $W_{i}=\left\langle w_{i1},w_{i2},w_{i3},\dots,w_{im}\right\rangle$, and compute the dominant shares with $ds_{i}=\max_{j}\left\{u_{ij}/w_{ij}\right\}$. In Parkes et al. (2015) study, weights are further defined to be normalised, such that $\sum_{i}w_{ir}=1$, for each resource $r$, and they use another "$\rho$" variable, at the construction of which the weights scale the normalised demands, to be used in their calculations. Again, it is notationwise involved and will not be elaborated on in more detail for the sake of simplicity.
@@
At this point, we believe that DRF and its derivatives offer a richer functionality than previously expressed. Even without weights, DRF is an implicit pricing mechanism, implicitly assigning value to resources on the basis of demand and supply dynamics. Inclusion of weights articulates it even stronger. For example, allowing users to purchase resources, or open-endedly invest in resources, and assigning their fraction of the total investment as the user’s weight, DRF can implement an implicit auctioning system. Further scrutinising this issue may be a research topic on its own, and is beyond the scope of the present study, thus we will suffice with pointing to it.
@@ 7 Conclusion
The allocation of multiple resources has become an increasingly important problem in the last decade and a half. The seminal study of Ghodsi and his colleagues has created a small literature of its own, since DRF brought many novelties and as much possibilities into this field. The study of Parkes and his colleagues contributes to this literature dearly, by providing a more rigorous theoretical framework to study the properties of DRF in higher resolution. The present study is one situated in this literature, offering yet another analysis method, and an improvement to the original algorithm to reach the same allocation in fewer steps.

### v2

@@ 7 Performance of PDRF
In the original DRF article it is pointed out that each allocation operation takes $\mathcal{O}(log(n))$ time, since the selection of lowest allocated dominant share may be done with a binary heap. However, the number of allocation operations is determined by the ratio of the available resource reserves to the demand averages. More rigorously, the expected execution time of the algorithm is given by:
@@
where, $D$ and $R$ denote the set of demands and resources, respectively, and $\mu_{dr}$ is the demand average for resource $r$. This can lead to performance problems in cases where this ratio is high.
The main contribution of PDRF is that in such cases it can reduce the time of execution dramatically, since it operates in linear time to the number of users and resources, i.e. $\mathcal{O}(m,n)$, and is not effected from this ratio.
PDRF executes $m\cdot n$ multiplications to find the fractional demands, $(m-1)\cdot n$ comparisons to find the $ds$ of each demand, n1n-1 comparisons to find $ds^{*}$, $n$ divisions for finding the ratio of $ds^{*}$ to each $ds$, $m\cdot n$ scaling of demands by the obtained ratios, $m\cdot(n-1)$ additions for summing up the scaled demands, $m$ divisions of resources by the demand sums, m1m-1 comparisons to find the minimum of the results of the divisions to decide the number of of cycle iterations, and $2\cdot m\cdot n$ operations for the final assignment of resources, all of which are linear time operations. It is rather straightforward that the heuristic method also takes $\mathcal{O}(n)$ steps.
@@ 8.1 Higher Order Cycles
Without loss of generality, we may consider the following single resource scenario where the demands are $2$, 44, and $10$, for users A, B and C, respectively, and the reserve is $35$ units. User C has the $ds^{*}$, so one cycle iteration consists of $1$ iteration for C, $2$ for B, $4$ for A, and it consumes $2\cdot 4+4\cdot 2+10\cdot 1=26$ units of the reserve. After one iteration of the cycle, $9$ units of resource is left and the cycle cannot take another turn. When we exclude User C, who has the $ds^{*}$, what we are left with is a secondary cycle, with now B having the new $ds^{*}$ and taking $1$ unit, and A taking $2$. The cycle takes one turn consuming 8 resources, and terminates.
@@ 8.2 Pareto Efficiency
The proof is based on the claim that ”progressive filling for DRF is equivalent to the algorithm presented in Figure 11111Since Figure $1$ is a scatter plot, we assume what is meant is Algorithm $1$. [1, p. 13]. Although it is indicated in the preceding paragraph that the progressive filling algorithm eliminates the saturated demands, no removal condition is given in Algorithm $1$, and neither it is mentioned in the sections and examples describing the algorithm. This might be a minor point eluding the authors’ attention or it might have been assumed that the selection process automatically excludes those demands, but the resulting algorithm as such, cannot achieve Pareto Efficiency in, though rare might they be, certain demand distributions and resource reserves.
At the moment we cannot provide an exact algebraic model for the construction of such configurations, but when the $ds^{*}$ is saturated, and there are more resources available for any other lower demand than its number of occurrences in the cycle, meaning, though saturated, $ds^{*}$ will come to be the least allocated task with the beginning of the next cycle, the algorithm terminates without allocating available resources to the other users.
An example we rather stumbled upon, which led to this analysis, consists of the same demand vectors in the original article, user A with the demand vector $\left\langle 1,4\right\rangle$ and user B with $\left\langle 3,1\right\rangle$, and the available capacities $\left\langle 59,19\right\rangle$. In such a setting, after 10 iterations, the algorithm halts allocating $\left\langle 2,8\right\rangle$ and $\left\langle 24,8\right\rangle$, respectively, and $\left\langle 33,3\right\rangle$ resources still available, which is sufficient for $3$ more allocations for user B, but since the least allocated dominant share belongs to A, and his demand is saturated, the algorithm cannot proceed. A short simulator script with this example can be found at the same github repository [10].
As in EDRF, this problem may easily be overcome by excluding the saturated demand from the demand set, but it further extends its execution time.
@@ 8.3 Weighting Policies
A weighting policy, assigns weights to users, and allocates the resources in proportion to these weights. Fairness, in turn, is defined over the allocations scaled by these weights. In [1], the question of weighting is addressed briefly, pointing to the fact that it suffices for such a generalisation to assign a weight vector to each user, $W_{i}=\left\langle w_{i1},w_{i2},w_{i3},\dots,w_{im}\right\rangle$, and compute the dominant shares with $ds_{i}=\max_{j}\left\{u_{ij}/w_{ij}\right\}$. In [4], weights are further defined to be normalised, such that $\sum_{i}w_{ir}=1$, for each resource $r$, and they use another $\rho$” variable, at the construction of which the weights scale the normalised demands, to be used in their calculations. Again, it is notationwise involved and will not be elaborated on in more detail for the sake of simplicity.
@@
At this point, we believe that DRF and its derivatives offer a richer functionality than previously expressed. Even without weights, DRF is an implicit pricing mechanism, implicitly assigning value to resources on the basis of demand and supply dynamics.
The inclusion of weights articulates this point even stronger. For example, allowing users to purchase weight tokens for resources, or open-endedly invest in weights for resources, and assigning their fraction of the total token sum or total investment as the user’s weight, DRF can implement an implicit auctioning system. Further scrutinising this issue may be a research topic on its own, and is beyond the scope of the present study, thus we will suffice with pointing to it.
@@ 9 Conclusion
The allocation of multiple resources has become an increasingly important problem in the last decade and a half. The seminal study of Ghodsi and his colleagues has created a small literature of its own, since DRF brought many novelties and as much possibilities into this field. The study of Parkes and his colleagues contributed to this literature dearly, by providing a more rigorous theoretical framework to study the properties of DRF in higher resolution.
The present study is one situated in this literature, offering yet another analysis method, and an improvement to the original algorithm to approximate the allocation at high accuracy in fewer steps.